a 25.00 ml sample of a weak acid is titrated with 0.225 M NaOH. a total of 12.20 ml of the NaOH is required to reach the equivalence point where the pH is 9.96. determine the value of pka for this weak acid
Answer #1

assume weak acid = HA

millimoles of NaOH added = 12.20 x 0.225 = 2.745

2.745 millimoles acid reacted

salt formed = 2.745

total volume = 25.00 + 12.20 = 37.2 mL

[salt] = 2.745 / 27.2 = 0.10 M

for salt of weak acid

pH = 1/2 [pKw + pKa + log C]

pH = 9.96

C = 0.10 M

9.96 = 1/2 [14 + pKa + log 0.10]

19.92 = 14 + pKa - 1  

19.92 = 13 + pKa

pKa = 6.92